OpenStudy (anonymous):

the product of radical 16 and radical 100

OpenStudy (amistre64):

\[\sqrt{n}*\sqrt{m}\to n^{1/2}m^{1/2}\to(nm)^{1/2}=\sqrt{nm}\]

OpenStudy (amistre64):

or 4*10